OpenStudy (anonymous):

[3c(4^2)-6c]/[5^2-4] simplify the expression completely

OpenStudy (debbieg):

\[\Large \frac{ 3c4^2-6c }{ 5^2-4 }\] right?

OpenStudy (anonymous):

right

OpenStudy (debbieg):

So... apply PEMDAS to the num'r and den'r (separately). Evaluate the exponents, then you'll have some multiplication in the num'r (the 3 * 4^2), then add/subtract the like terms. You'll finally get to where you have a single term (a monomial) in the num'r, and just a number in the den'r. You might be able to reduce at that point. ;)

OpenStudy (anonymous):

i can't understand please the answer?

OpenStudy (debbieg):

As is the site policy, I will not give you the answer. http://openstudy.com/code-of-conduct Anyway, it is of no use to you to have the answer without understanding how to do the problem, as clearly your teacher thinks you should be able to do. If you want me to help you UNDERSTAND HOW to do it, I will. If you don't want to understand it, then I won't bother. Up to you.

OpenStudy (anonymous):

can you draw me how to do it

OpenStudy (anonymous):

please

OpenStudy (debbieg):

What is 4^2? What is 5^2?

OpenStudy (anonymous):

=16 =25

OpenStudy (debbieg):

Ok good... so now we have \[\Large \frac{ 3c(16)-6c }{ 25-4 }\] Now do the product in the first term of the num'r: 3*c*16 = ?

OpenStudy (anonymous):

c=16/3 c=5.3

OpenStudy (debbieg):

Ack... no. First of all, we aren't SOLVING for c. it's a TERM not an EQUATION. Just simplify the product: 3*c*16 = ? E.g., 2*m*12 = 24m 5*z*(-7)=-35z so what is 3*c*16 = ?

OpenStudy (anonymous):

=48

OpenStudy (debbieg):

almost.... don't forget your c! So you get 48c for that term. Now we have:

OpenStudy (anonymous):

48c-6c

OpenStudy (debbieg):

\[\Large \frac{ 48c-6c }{ 25-4 }\] OK, now one at a time we'll do the subtraction. In the den'r, you have 25 - 4 =? In the num'r, you have 48c - 6c =? (remember that you do this kind of addition/subtraction just by adding/subtracting the coefficients).

OpenStudy (anonymous):

42c/21

OpenStudy (debbieg):

Excellent! And now you can reduce..... to end up with ?

OpenStudy (anonymous):

14c/7

OpenStudy (debbieg):

ok... and one more - you can still reduce further, you still have a common factor. :)

OpenStudy (anonymous):

i dont know

OpenStudy (debbieg):

Hmmm...... well, 14 = ? * 7 ??

OpenStudy (anonymous):

2*7=14

OpenStudy (debbieg):

Right, so 14/7 = ?

OpenStudy (debbieg):

It's not fully simplified until you have cancelled all common factors.... so far you have: \(\Large \dfrac{ 48c }{ 21} =\dfrac{ 14c }{ 7} =\dfrac{ ?c }{ ?} \)

OpenStudy (anonymous):

7/7

OpenStudy (debbieg):

oops, that should be 42c/21

OpenStudy (debbieg):

14/7 = 7/7 ??

OpenStudy (debbieg):

does 14 = 7 ?? :) C'mon, you know how to reduce this....

OpenStudy (debbieg):

\(\Large \dfrac{ 42c }{ 21} =\dfrac{ 14c }{ 7} =\dfrac{ ?c }{ ?} \)

OpenStudy (anonymous):

7c/7

OpenStudy (debbieg):

\(\Large \dfrac{ 14 }{ 7} \neq\dfrac{ 7 }{ 7} \)

OpenStudy (debbieg):

You told me that 2*7 = 14 So \(\Large \dfrac{ 14c }{ 7} =\dfrac{7\cdot2c }{ 7} \) right??

OpenStudy (anonymous):

i dont know what to do

OpenStudy (debbieg):

If you don't mind my asking, what class is this for? What grade are you in? (I'm just wondering... it helps me understand what your math knowledge is.)

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

im 9th its algebra 1

OpenStudy (debbieg):

\(\Large \dfrac{ 14c }{ 7} =\dfrac{7\cdot2c }{ 7}=\dfrac{7}{ 7}\cdot2c\)

OpenStudy (debbieg):

\(\Large =\dfrac{\cancel7}{ \cancel7}\cdot2c=1\cdot2c\)

OpenStudy (anonymous):

1.2c is the answer?

OpenStudy (anonymous):

or 1*2c

OpenStudy (debbieg):

OK, you should probably review your multiplication tables and basic products, factors, etc. it isn't 1.2c.... it's 1 TIMES 2c. You don't need the 1, it was just to show you why the factors reduce as they do. \(\Large =\dfrac{\cancel7}{ \cancel7}\cdot2c=1\cdot2c=2c\)

OpenStudy (debbieg):

But it is VERY IMPORTANT that you UNDERSTAND how to do this. I walked you through it, but I can't help you on your test. :) YOU need to grasp it.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so much thanks

OpenStudy (debbieg):

You're welcome. :) happy to help!